OpenStudy (anonymous):

What is f(x) = x2 – 3x + 2 for f(-1)? -2 0 4 6

OpenStudy (anonymous):

plug -1 in for x (-1)2 - 3(-1) + 2 = -2 +3 + 2 = 3

OpenStudy (anonymous):

f(-1) just means to put -1 everywhere you have x

OpenStudy (anonymous):

@Sheng That answer is wrong.

OpenStudy (anonymous):

im interpreting x2 = 2x

OpenStudy (anonymous):

if it's x^2, it's missing a "^"

OpenStudy (anonymous):

so is it A?

OpenStudy (anonymous):

assuming x^2, it'll be 6

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

@Sheng It is not customary to just give out the answers.